OpenStudy (anonymous):

simplify. 2.7 x 10^-8 over 9 x 10^-4

OpenStudy (kropot72):

\[\frac{2.7\times 10^{-8}}{9\times 10^{-4}}\] To simplify, first multiply the numerator and the denominator by 10^4. After doing that, divide the numerator and the denominator by 9.

OpenStudy (zehanz):

Or try this: \[\frac{ 2.7 \times 10^{-8} }{ 9 \times 10^{-4} }=\frac{ 27 \times 10^{-9 }}{ 9 \times 10^{-4} }=\frac{ 27 }{ 9 } \times \frac{ 10^{-9} }{ 10^{_-4} }= ...\]Work out the two quotients and for the right one use the rule:\[\frac{ a^b }{ a^c }=a^{b-c}\]